# Retrieval-Augmented Generation for AI-Generated Content: A Survey

### v3

@@ IV-B6 Others
There are several other code-related tasks that adopt RAG paradigm. In  [288] for numerical reasoning task, the Chain-of-Thought is replaced with the programs as the intermediate reasoning step, and dense retrieval-based similar examples are augmented in prompt for downstream LLMs. De-fine [289] tries to resolve intricate tasks using programs. It follows the paradigm in SKCODER, retrieves relevant pairs of query and code, then produces sketch template for real program generation. After generation, it combines the feedback to refine the answer with the same generator. The refined programs, regarded as optimal results, are added back to the codebase for subsequent serving. E&V [290] leverages an LLM-based agent for program static analysis. The agent uses source code retrieval, pseudo-code execution, execution specifications verification, and other tools to form intermediate results. The retrieval is implemented by AST pattern matching. Code4UIE [291] leverages code representation for information extraction tasks. It retrieves relevant examples through dense retrieval, and constructs in-context learning prompt using the retrieved queries and their corresponding codes. StackSpotAI [292] builds an AI coding assistant, which incorporates many code-related tasks. It implements an RAG component, identifying the most relevant pieces of information which serve as the context for GPT generation. InputBlaster [293] aims to generate unusual text input that could cause mobile app crash. It combines retrieved relevant buggy text with valid input to form the prompt for generation.
@@ IV-C1 Knowledge Base Question Answering
Query-based RAG is the mainstream approach. For a given query, Unseen Entity Handling [53] retrieves topic entities through FreeBase [294], and concatenates the query with the entity for an encoder-decoder to generate SPARQL output. CBR-KBQA [54] retrieves relevant questions and corresponding logical form answers with roberta-based deep retrieval, then concatenates the question and the retrieved pairs for encoder-decoder transformer model. It also revises the final generation result to align the generated relations with relations present in the local neighborhood of the query entity in the knowledge graph. GMT-KBQA [52] first retrieves relevant entities and relations through bert-based deep retrieval, then concatenates the information for T5 generation. To improve the retrieval result, it leverages cross-encoder to rerank the candidates, and uses the same T5 structure to conduct relation classification and entity dismbiguation. RNG-KBQA [123] enumerates candidate logical forms in the knowledge graph, then ranks the candidates and concatenates them with query to form the prompt input to generate final logical form through a T5 model. Based on this idea, TIARA [124] also retrieve entity and schema besides logical forms, while a following work [295] retrieves top-k questions with BM25. Uni-Parser [133] retrieves relevant entities from knowledge graph using mention detection, cross-encoder ranker, and 2-hop paths extraction. It also considers enumerating tables and columns from databases. On obtaining the relevant information, it concatenates the top-k primitives with the query and generates logical forms through T5. BLLM augmentation [135] uses TIARA as the retrieval for relevant knowledge base elements, then performs in-context learning through black-box LLM such as GPT-4 for generating logical forms. ECBRF [134] follows the case-based reasoning paradigm [296], retrieving similar triplet with dense retriever and constructing prompt input for BART or GPT-2 in-context learning. FC-KBQA [297] extracts relevant class, relation, and entity given a question. For class and relation, it uses BM25 as retriever and a Bert-based cross-encoder as re-ranker. For entity, it follows the mention detection paradigm. To compose all the component candidates, it applies T5 model to generate logical expression. StructGPT [298] extracts relevant information, including triplets and nearest entities, to form prompt for subsequent LLM. KAPING [299] builds prompt including the user query and the retrieved relevant facts (through entity matching), then generates the answer through LLM. Another work [300] also follows the retrieve-then-generate paradigm, replacing the retrieval with a relation distribution generation model for weighted triplets. Retrieve-Rewrite-Answer [301] first retrieves subgraph using hop prediction, relation path prediction, and triplet sampling. It then performs KG-to-text and zero-shot generation with retrieved subgraphs as prompt. Keqing [302] first decomposes a complext question into simple sub-questions through finetuned LLM, then retrieves similar sub-question template by dense retriever RoBERTa to extract candidate entities from knowledge graph, and finally generates answer through ChatGPT given relevant entities as context input. To probe the deep understanding of natural language in LLMs with formal languages, a research work [303] explores the capability of formal language understanding and formal language generation. It leverages retrieved pairs to perform in-context learning. For understanding, it uses tree edit distance to retrieve similar logical forms, while for generation, it uses BM25 to retrieve similar natural language queries. Interactive-KBQA [304] treats LLM as agent and KG as environment. In each step, the LLM conducts entity-linking and one-hop retrieval on KG, then generates current thought and action, until obtaining the final answer.
@@
Latent representation-based RAG is also employed in KBQA. ReTraCk [305] links entities using mention detection, and retrieves schema items using dense retriever Bert. It then generates logical forms by LSTM, incorporating retrieved items through knowledge-specific rules. SKP [145] concatenates triplets with the query and uses fusion-in-decoder technique in inference. It adds a pretraining stage with a knowledge-aware MLM loss on triplets and knowledge constrastive loss with respect to the retrieved items. DECAF [144] forms Resource Description Format knowledge base triplets into sentences, then concatenates sentences with the same head entity into documents. It retrieves relevant documents using BM25 sparse retrieval or Bert dense retrieval, then leverages Fusion-in-Decoder technique to generate logical form and direct answer given each (query, document) pair as input. It combines the output to obtain the final answer. KD-CoT [146] uses the same dense retriever and fusion-in-decoder generator as DECAF. It follows a Chain-of-Thought paradigm, iteratively performing retrieval, generation, and verification until the CoT is finished.
@@ IV-C2 Knowledge-augmented Open-domain Question Answering
The use of latent representation-based RAG, particularly the fusion-in-decoder technique, is prevalent for knowledge-augmented open-domain question answering. UniK-QA [143] concatenates the text forms of the components in a triplet and build document pool for retrieval. For a given question, it leverages dense retriever for relevant documents, then performs fusion-in-decoder technique to incorporate the information for answer generation.
@@
KG-FiD [306] conducts the retrieval and generation as normal FiD. It add re-ranking in two ways: the first is to use a graph attention network on the graph formed by retrieved documents; the second is to use the hidden states in the generator. OREOLM [307] empowers LLM with knowledge reasoning paths. Concretely, it uses entity linking to determine the initial state, then conducts contextualized random walk on KG to get reasoning paths, whose entity value memory are combined into the hidden states of LLM for better generation. GRAPE [308] constructs bipartite graph for each pair of question and retrieved passage, then builds bipartite graph on entity for fusing knowledge. It leverages FiD as backbone model and generate answers. SKURG [309] forms a knowledge graph using text and image data, then updates each source’s representation with their relevant sources. It then uses a specially designed decoder to iteratively retrieve and generate. It conducts cross-attention with all the sources, then retrieves the source with highest score and concatenates to the original input embedding; if a gate score does not exceed the threshold, it starts to generate the real answer, otherwise the retrieval stage re-starts.
With the rapid development of LLMs, query-based RAG is emerging as a new standard. DIVKNOWQA [310] develops a retrieval-augmentation tool, including sparse retrieval on structured knowledge, dense retrieval on texts, and sparql generation on KB. Through CoT-based LLM, it retrieves and re-ranks in each step, and generates the final answer. KnowledGPT [311] uses generated code to retrieve from both public and personal knowledge bases, so as to build prompt for LLM question answering. EFSUM [312] generates evidence-focused summary with retrieved relevant facts, then optimizes the summary to align the QA-specific preference with another generator and the filters for helpfulness and faithfulness. GenTKGQA [313] conducts subgraph retrieval through relation ranking and time mining, then employs GNN to incorporate structural and temporal information into virtual token representations for subsequent LLM generation. KnowledgeNavigator [314] analyzes complex questions and performs retrieval on knowledge graph through iterative filtering of relations with respect to core entities, so as to obtain relevant triplets. It then includes the triplets into prompt for generation.
@@ IV-C3 Table for Question Answering
Fusion-in-decoder [35] style RAG is often used for table QA. EfficientQA [315], a competition held in NeurIPS 2020, witnessed the proposal of numerous retrieval-reader systems that rely on textual and tabular data. Dual Reader-Parser [316] re-ranks the retrieved textual and tabular data for generation. Convinse [317] retrieves information from heterogeneous resources (including knowledge bases, tables, and texts) after question understanding. CORE [318] retrieves relevant tables and passages through dense representation, then re-ranks the retrieved results using query-generation score from a T0 linker. It uses FiD for final generation. RINK [319] follows the retriever-reranker-reader paradigm for table-augmented question answering. It designs a set-level reader-inherited reranker to get the relevance score of blocks (table segments). TAG-QA [320] retrieves from both tables and texts: for tables, it converts tables to graphs then performs GNN to select relevant contents; for texts, it uses BM25 for retrieval. It then leverages FiD for answer generation.
Tables can be incorporated into prompts for query-based RAG. T-RAG [321] retrieves relevant tables given a user query, then concatenates the query with the tables as a prompt to generate the answer through BART. OmniTab [322] conducts multi-task training to improve the performance of table question answering model. It retrieves relevant tables given a query, then concatenates them for masked-token pre-training. CARP [323] retrieves relevant tables and passages using entity linking, then extracts hybrid chain of retrieved knowledge, which is later used to construct prompt for generation. StructGPT [298] extracts relevant information from knowledge graph, table, or database, to form prompt and generate answers through LLMs. cTBLS [324] retrieves relevant tables through dense retrieval, then for each query, it forms prompt with ranked tables for answer generation. A recent work [325] first uses table-to-text techniques to integrate tabular data into corpora, then conducts experiments on both finetuning and RAG for question answering.
@@ IV-C4 Others
Prototype-KRG [326] retrieves knowledge facts and dialogue prototypes, then integrates them into the GRU-based generator by both hidden states and logits. SURGE [327] retrieves relevant subgraphs using dense retrieval, then adds them into the input of the generator for dialogue generation. RHO [328] fuses KG embedding of relevant entities and relations into textual embeddings during the generation of open-domain dialogue. K-LaMP [329] retrieves entity in history queries to construct prompt for query suggestion. ReSKGC [147] linearizes all training triplets into text by concatenation, then retrieves relevant triplets using BM25, and generates completed triplet using T5 with fusion-in-decoder. G-Retriever [330] retrieves relevant nodes and edges from graph-based data, then constructs subgraph and performs graph prompt tuning for question answering based on textual graphs.
@@ IV-D1 Image Generation
The retrieval process can not only help yield high-quality images even for rare or unseen subjects, but also reduces the parameter count and computational expense [45, 137, 149, 150, 148, 331, 151, 152]. For GAN-based model, RetrieveGAN [45] employs a differentiable retrieval process to select compatible image patches for generation, with Gumbel-softmax trick and end-to-end training. IC-GAN [137] models data as a mix of conditional distributions around each training instance, conditioning both the generator and discriminator on these instances, and can control the semantics and style by swapping class labels or conditional instances. Recently, diffusion models beat GANs on image generation [332]. KNN-Diffusion [149] trains a text-to-image diffusion model without text data, by conditions the model on CLIP joint embedding of the instance and $k$-nearest neighbors from image database. These $k$-NN embeddings bridge the text-image distribution gap and allow image generation from different domains by swapping the database. Similarly, RDM [150] conditions diffusion or autoregressive models on CLIP embeddings of external image databases. It enables post-hoc conditioning on class labels, text prompts and zero-shot stylization [151]. Beyond retrieving only images, Re-imagen [148] conditions on both text prompts and retrieved image-text pairs for text-to-image generation. Interleaved classifier-free guidance is also proposed to balance the alignment between text prompts and retrieval conditions. To avoid information loss of CLIP embeddings and access to all visual condition, Retrieve&Fuse [331] concatenates the retrieved conditional image and the noised image before each attention block in U-Net, and allowing interaction via self-attention. RPG [79] retrieves representative images to construct informative in-context examples (i.e., image-region pairs), and utilizes multi-modal chain-of-thought reasoning [333] to plan out complementary subregions for compositional text-to-image diffusion.
@@ IV-D2 Image Captioning
Retrieval-augmented image captioning typically synthesises description with a collection of retrieved captions, instead depending only on the input image. MA [164] augments via a memory bank, built with historical context and target word of image-text training set, and queried during inference with the current context. Vocabulary distribution is computed based on retrieved entries, and interpolated with the original prediction. In adversarial training, RAMP [334] employs retrieved captions as reference for discriminator training, prompting the generator to make full use of retrieved captions. The memory-augmented attention and copying mechanism are also exploited to better use. The RA-Transformer [46] and EXTRA [335], both retrieval-augmented transformer-based captioning models, utilize cross-attention over encoded retrieved captions.
@@
EXTRA, as depicted in Fig. 10, jointly process the image and retrieved captions with V&L encoder, such that the decoder attends to both visual and linguistic contexts. Beyond retrieved captions, REVEAL [336] uniformly encodes and retrieves multi-modal world knowledge, including image-text pairs, question answering pairs, and knowledge graph triplets, which is then integrated with image features by retrieval score-aware attention module. Straightforwardly, SMALLCAP [47] employs a CLIP vision encoder and a LLM decoder, linked by trainable cross-attention layers, where retrieved captions serve as input-specific in-context examples for prompt for enhancement. For remote sensing image, CRSR [337] enhances retrieved captions with semantic refinement, i.e. filters out misleading details and emphasizes visually salient content. Besides, the visual features is also enriched by transformer network with learnable queries, capturing more intricate details within the images
@@ IV-D3 Others
There also exist many retrieval augmented works for other image-related tasks. For visual question answering (VQA), PICa [338] leverages GPT-3’s implicit knowledge retrieval and reasoning capabilities, which converts images into textual descriptions, then prompts GPT-3 to predict answers based on these descriptions and the question, finally ensembles multi-query results. RA-VQA [339] identifies a limitation that the retrieval in previous work is trained separately from answer generation, and propose a joint training scheme that integrates differentiable retrieval with answer generation, enabling end-to-end training. For visually grounded dialogue, KNN-based Information Fetching (KIF) [340] enhances generative Transformer for dialog modeling. Each KIF module learns a read operation to access fixed external knowledge. Maria [341], a neural conversation agent, enhances dialog generation by leveraging visual experiences retrieved from a large-scale image index as extra context. For multi-modal machine translation, which aims to improve NMT with multi-modal information, [342] incorporates visual information at the phrase level to address the sparsity of paired sentence-image, employing a conditional VAE to filters out redundant visual information from sentence-image datasets.
@@ IV-E1 Video Captioning
Video captioning is to describe the visual content with a descriptive utterance. KaVD [343] generates news video caption with background knowledge mined from topically related documents such as named entities and events. R-ConvED [48] introduces retrieval-augmented mechanism to facilitate the word prediction. It uses Dual Encoding [109] for video-text retrieval, and proposes a convolutional encoder-decoder network for generation. For a given input video, R-ConvED first retrieves top-k relevant sentences and their corresponding video from training set, then feeds these pairs and the input video into the generator separately. The obtained decoder hidden states are combined through attention-like read operation, so that the target word can be predicted using the final representation. CARE [160] utilizes visual encoder, audio encoder, and text encoder for frame, audio, and retrieved texts, respectively. It uses CLIP as retriever, and transformer decoder as generator. The embeddings of the three modalities are combined to augment the decoder, producing global semantic guidance which attends the input embedding, and local semantic guidance which attends the attention layer. EgoInstructor [49] generates captions for first-person view videos. It retrieves relevant exocentric videos and corresponding texts via dense retrieval, then encodes the input egocentric video and the retrieved exocentric videos through a CLIP-based visual encoder and a transformer-decoder-based bottleneck module. Then it generates captions through decoder-based LLM which takes the retrieved texts as input and interacts with encoded videos in gated cross-attention layer.
@@ IV-E2 Video QA&Dialogue
Video QA&Dialogue generates single or multiple-round responses in alignment with video content. For video question answering (VideoQA), MA-DRNN [344] leverages the differentiable neural computer (DNC) with an external memory, for storing and retrieving useful information in questions and videos, and modeling the long-term visual-textual dependence. Given the video input, R2A [345] retrieves semantically similar texts by multi-modal model, e.g. CLIP, and query LLM with both the question and the retrieved texts. For video-grounded dialogue, [346] proposes TVQA+ dataset which enables to retrieve relevant moments and visual concepts to answer questions about videos, and also proposes Spatio-Temporal Answerer with Grounded Evidence (STAGE) to exploit it. VGNMN [347] also extracts visual cues from videos, while the retrieval is carried out using neural module networks parameterized by entities and actions in previous dialogues.

### v4

@@ IV-B6 Others
There are several other code-related tasks that adopt query-based RAG paradigm, incorporating similar examples to construct prompts. Jie et al. [242] used programs as the intermediate step in numerical reasoning. De-fine [243] uses programs to solve complex tasks. It refines the answer generated by query-based RAG, then adds the refined programs back to the retrieval source. For program static analysis, E&V [244] leverages an LLM agent to form intermediate results with AST-based source code retrieval, pseudo-code execution, execution specifications verification, and other tools. Code4UIE [245] performs information extraction through code representation. StackSpotAI [246] builds an AI coding assistant with an RAG component. InputBlaster [247] generates unusual text input that could cause mobile app crash.
@@ IV-C1 Knowledge Base Question Answering
Query-based RAG is the mainstream approach. Unseen Entity Handling [53] uses FreeBase [248] to retrieve topic entities, which are combined with query to generate SPARQL output. CBR-KBQA [54] combines the query and the retrieved (query, logical form) pairs for generation. It also revises the final result to align with the relations present in the knowledge graph. GMT-KBQA [52] re-ranks the retrieved entities and relations, and conducts relation classification and entity disambiguation before generation. RNG-KBQA [81], TIARA [82], BLLM augmentation [249], and Shu et al. [250] re-rank the candidate logical forms or entities from the knowledge graph for prompt construction. Uni-Parser [91] includes entities from mention detection, 2-hop paths extraction, and tables from databases into generator input. ECBRF [92] follows the case-based reasoning paradigm [251], retrieving similar triplet to build prompt input. FC-KBQA [252] extracts relevant classes, relations, and entities from BM25 or mention detection, StructGPT [253] extracts relevant triplets and nearest entities, and KAPING [254] extracts relevant facts through entity matching. Sen et al. [255] replaced the retrieval with a relation distribution generation model for weighted triplets. Retrieve-Rewrite-Answer [256] retrieves subgraphs into prompts using hop prediction, relation path prediction, and triplet sampling. Keqing [257] decomposes a complex question into simple sub-questions through LLM, then retrieves sub-question templates and extract candidate entities from knowledge graph, and finally generates the answer through ChatGPT. Liu et al. [258] leveraged retrieved pairs to explore the capability of formal language understanding and generation. Interactive-KBQA [259] employs the LLM as an agent, which conducts entity-linking on KG and generates current thought and action until obtaining the final answer.
@@
Latent representation-based RAG is also employed for KBQA. ReTraCk [260] retrieves entities and schemas through mention detection and dense retrieval. It generates logical forms using LSTM, using retrieved items through knowledge-specific rules. SKP [109], DECAF [108], and KD-CoT [110] all retrieve triplets and conduct fusion-in-decoder [35] RAG. KD-CoT also follows a chain-of-thought paradigm, iteratively performing retrieval, generation, and verification.
@@ IV-C2 Knowledge-augmented Open-domain Question Answering
Latent representation-based RAG, especially the fusion-in-decoder [35] technique, is prevalent for knowledge-augmented ODQA. UniK-QA [107], KG-FiD [261], GRAPE [262] all apply the fusion-in-decoder technique. They incorporate triplet-based documents, re-ranked documents through KG, and bipartite graph for pairs of question and passage, respectively.
@@
OREOLM [263] empowers LLM with knowledge reasoning paths, integrating the entity value memory derived from contextualized random walk paths on KG into the hidden states of the LLM. SKURG [264] performs iterative retrieval and generation, using cross-attention to incorporate data sources into the input embedding. It uses a gate score to determine whether to re-start retrieval or to generate the real answer.
With the rapid development of LLMs, query-based RAG is emerging as a new standard. DIVKNOWQA [265] retrieves from multiple sources using different techniques. It iteratively retrieves and re-ranks the data before generating the final answer. KnowledGPT [266] uses generated code to retrieve from both public and personal knowledge bases. EFSUM [267] optimizes the evidence-focused summary after facts-augmented generation, so as to align the QA-specific preference for helpfulness and faithfulness. GenTKGQA [268] employs GNN (graph neural network) to integrate structural and temporal information from subgraph retrieval into virtual token representations. KnowledgeNavigator [269] performs retrieval on KG through iterative filtering of relations with respect to core entities, so as to obtain relevant triplets.
@@ IV-C3 Table for Question Answering
Fusion-in-decoder [35] style RAG is often used for table QA. EfficientQA [270], a competition held in NeurIPS 2020, witnessed the proposal of numerous retrieval-reader systems that rely on textual and tabular data. Dual Reader-Parser [271] and CORE [272] both re-rank the retrieved textual and tabular data for generation. Convinse [273] retrieves information from knowledge bases, tables, and texts after question understanding. RINK [274] designs a set-level reader-inherited re-ranker to get the relevance score of table segments. TAG-QA [275] retrieves tables and texts through GNN (after table-to-graph conversion) and BM25, respectively.
Tables can be integrated into prompts for query-based RAG. Both T-RAG [276] and OmniTab [277] concatenates the retrieved tables with the query to generate the answer. CARP [278] extracts hybrid chain of retrieved tables and passages for prompt construction. StructGPT [253] retrieves from multiple sources including KGs, tables, and databases. cTBLS [279] forms prompts with ranked tables after retrieval. Min et al. [280] integrated tabular data through table-to-text techniques, then experiments on both finetuning and RAG.
@@ IV-C4 Others
Prototype-KRG [281] integrates retrieved knowledge facts and dialogue prototypes into a GRU model through both hidden states and logits. SURGE [282] combines relevant subgraphs into the input for dialogue generation. RHO [283] fuses KG embedding of relevant entities and relations into textual embeddings during dialogue generation. K-LaMP [284] retrieves entities in history queries to construct prompt for query suggestion. ReSKGC [111] retrieves relevant triplets to complete triplet using Fid. G-Retriever [285] retrieves nodes and edges from textual graphs to construct subgraph and perform graph prompt tuning for QA.
@@ IV-D1 Image Generation
The retrieval process can not only help yield high-quality images even for rare or unseen subjects, but also reduces the parameter count and computational expense [45, 94, 103, 104, 102, 286, 105, 106]. For GAN-based model, RetrieveGAN [45] uses a differentiable retriever for image patch selection, facilitating end-to-end training. IC-GAN [94] models data as conditional distributions around each training instance, conditioning both the generator and discriminator on these instances. Recently, diffusion models beat GANs on image generation [287]. KNN-Diffusion [103] and RDM [104] train diffusion models conditioned on CLIP embeddings and image neighbors, enabling post-hoc conditioning on labels, prompts, and zero-shot stylization [105]. Beyond only images, Re-imagen [102] extends retrieval to image-text pairs for text-to-image generation, with interleaved guidance to balance the alignment between prompts and retrieval conditions. Retrieve&Fuse [286] prevents information loss of CLIP embeddings by concatenating retrieved and noised images before each U-Net attention block, allowing fully interaction via self-attention. RPG [288] retrieves representative images to construct in-context examples, and utilizes chain-of-thought reasoning [289] to plan out complementary subregions for compositional text-to-image diffusion.
@@ IV-D2 Image Captioning
Retrieval-augmented image captioning typically synthesises description with a collection of retrieved captions. MA [120] augments via a memory bank, built with historical context and target word of image-text training set, and queried with inference context. Vocabulary distribution is computed based on retrieved entries, and interpolated with the original prediction. In adversarial training, RAMP [290] takes retrieved captions as discriminator reference, and employs memory-augmented attention and copying mechanisms for better utilization of retrieved captions.
@@
The RA-Transformer [46] and EXTRA [291], both retrieval-augmented transformer-based captioning models, utilize cross-attention over encoded retrieved captions. Beyond caption retrieval, REVEAL [292] uniformly encodes and retrieves multi-modal world knowledge, integrated with retrieval score-aware attention. Directly, SMALLCAP [47] employs a CLIP vision encoder and a LLM decoder, with retrieved captions serving as input-specific in-context examples. For remote sensing images, CRSR [293] refines retrieved captions, filtering out misleading details and emphasizing visually salient content.
@@ IV-D3 Others
There also exist many retrieval augmented works for other image-related tasks. For visual question answering (VQA), PICa [294] converts images into textual descriptions, prompts GPT-3 and ensembles multi-query results. RA-VQA [295] enables an end-to-end training with differentiable retrieval for answer generation. For visually grounded dialogue, KIF [296] and Maria [297] enhances dialog generation with external knowledge like visual experiences. In multi-modal machine translation, [298] incorporates visual information at the phrase level to improve NMT with multi-modal information.
@@ IV-E1 Video Captioning
Video captioning translates the visual content into descriptive utterances. KaVD [299] generates news video caption with background knowledge in related documents like named entities and events. R-ConvED [48] retrieves relevant sentences and videos via Dual Encoding [69], and predicts the target word with a convolutional encoder-decoder network. CARE [116] combines three modalities data, i.e. frame, audio, and retrieved texts, to provide both global and local semantic guidance as augmentation. EgoInstructor [49] focuses on first-person videos, retrieves relevant exocentric videos and texts, and generates captions through LLM via cross-attention with encoded videos
@@ IV-E2 Video QA&Dialogue
Video QA&Dialogue generates single or multiple-round responses in alignment with video content. For VideoQA, MA-DRNN [300] stores and retrieves useful information in queries and videos with external memory, therefore models the long-term visual-textual dependence. R2A [301] retrieves semantically similar texts by CLIP, and prompts LLM with both the query and the retrieved texts. For Video Dialogue, [302] proposes TVQA+ dataset to enable relevant moments and visual concepts retrieval, and designs corresponding spatio-temporal-aware generator. VGNMN [303] extracts visual cues from videos, while the retrieval process is parameterized by entities and actions in previous dialogues.